# Evolution of flat bands in MoSe2/WSe2 moiré lattices: A study combining machine learning and band unfolding methods

### v1

@@ Abstract
Moiré lattices have served as the ideal quantum simulation platform for exploring novel physics due to the flat electronic bands resulting from the long wavelength moiré potentials. However, the large sizes of this type of system challenge the first-principles methods for full calculations of their electronic structures, thus bringing difficulties in understanding the nature and evolution of the flat bands. In this study, we investigate the electronic structures of moiré patterns of MoSe2/WSe2 by combining ab initio and machine learning methods. We find that a flat band with a bandwidth of about 5 meV emerges below the valence band edge at the K point for the H-stacking at a twist angle of 3.89. Then, it shifts dramatically as the twist angle decreases and becomes about 20 meV higher than the valence band maximum for the twist angle of 3.15. Multiple ultra-flat bands emerge as the twist angle is reduced to 1.7. Band unfolding reveals that the ultra-flat bands formed by the $\Gamma$ and K valleys show distinct behaviors. The $\Gamma$-valley flat bands are sensitive to the interlayer coupling, thus experiencing dramatic changes as the twist angle decreases. In contrast, the K-valley flat band, which shows a weak dependence on the interlayer coupling, is mainly modulated by the long wavelength moiré potential. Therefore, a relatively small angle (2.13) is required to generate the K-valley flat band, which experiences a transition from the honeycomb to the triangular lattice as the twist angle decreases.
@@ I INTRODUCTION
The tbTMDs can exhibit much richer geometric and electronic structures than tBLG [50, 51]. For instance, they can be obtained by twisting the TMD bilayers either in R-stacking or in H-stacking. Namely, their structure is dependent on the relative orientation of the two layers. As a result, they show interesting geometric patterns such as honeycomb and triangular domain wall networks depending on the stacking order. In addition, the TMD bilayers possess two different groups of valleys, i.e., the $\Gamma$ and K valleys. The flat bands generated by the two valleys can be used to model different physics. The $\Gamma$-valley flat bands were proposed to realize honeycomb and kagome lattice models [47], while the K-valley flat bands can be used to study the triangular Hubbard model [40, 41, 52, 53]. Moreover, the two valleys are energetically close. This feature along with the stacking order and twist angle can give rise to complicated electronic structures for tbTMDs. For instance, for the twisted bilayer of WSe2 with an angle of 5.1 (a twist away from R-stacking) only the $\Gamma$-valley flat bands were observed [54, 55]. As the twist angle decreases to 3, both the $\Gamma$- and K-valley flat bands were observed [45]. For the twists near H-stacking, $\Gamma$-valley flat bands were observed for WSe2 at angles of 57.4 and 57.5 [56, 45]. When the angle is slightly increased to 57.72, the K-valley flat band appears and becomes 129 meV higher than the $\Gamma$-valley one [57].
@@ II METHODS
We use a DP generated by the DeePMD-kit package to model the interatomic interactions for the MoSe2/WSe2 moiré lattices. The training data were obtained by DFT calculations using the Vienna Ab-initio Simulation Package (VASP) [63]. The projector augmented wave (PAW) method was used to construct pseudopotentials [64]. The energy cutoff for plane waves was 270 eV and the electronic exchange-correlation functional was parametrized by the formalism proposed by Perdew, Burke, and Ernzerhof (PBE) within the generalized gradient approximation [65]. The 2D Brillouin zone (BZ) was sampled using a 5$\times$5 $k$-mesh. The van der Waals forces between layers were corrected using the DFT-D3 method [66]. Then, the LAMMPS program incorporating the DP model was used for structural relaxations [67, 68, 69], for which the threshold force is 10-6 eV/Å. The band structures were obtained using the DeepH-pack package. The atomic-orbital-based $ab$ $initio$ computation program at UStc (ABACUS) [70, 71] was used to generate the training data for DeepH. The pseudo-atomic orbitals used in the ABACUS calculations are Se-$2s2p1d$, Mo-$4s2p2d1f$, and W-$4s2p2d2f$ with a radius cutoff of 10 a.u. and an energy cutoff of 100 Ry. Unfolded bands were obtained using the KPROJ program [72], which projects the wave function of the moiré lattices onto the $k$-points in the BZ of the primitive cell.
@@ III.1 DP and DeepH models
For the DP model, the dataset for training was generated by $NVT$ $ab$ $initio$ molecular dynamic (AIMD) simulations using VASP for a few structures, which include 2 $\times$ 2 supercells of the high-symmetry stacking orders shown in Fig. 1, those obtained by a relative shifting of the top layer in the stacking orders, and moiré structures with an angle of 21.79, i.e., $H$-$21.79$ and $R$-$21.79$. In addition, these structures under biaxial strains in the range from -2.5% to 2.5% were also included. For the DeepH model, structures for generating the dataset are 3 $\times$ 3 supercells of the high-symmetry stacking orders, the shifted ones, and the twisted structures with an angle of 13.17. The Hamiltonian hopping and overlapping matrix elements were obtained using ABACUS. Benchmark tests for the DP and DeepH models are shown in Fig. 2. It should be mentioned that we also calculated phonon dispersions for $H$-$13.7$ [Fig. 2(f)] and $R$-$13.7$ [Fig. 2(h)] using the DP model interfaced with LAMMPS and calculated electronic bands for $H$-$21.79$ [Fig. 2(j)] and $R$-$21.79$ [Fig. 2(l)] using DeepH, of which the structures were not included in the training. However, the results from the deep-learning models show good agreements with those from DFT calculations, suggesting the high precision of the models.
@@ III.2 Geometric structures
Figure 3(a) shows a rigid moiré structure for $H$-1.7, which exhibits a triangular pattern. The pattern is then changed into a hexagonal domain-wall network after structural relaxation due to geometric reconstruction [see Fig. 3(b)]. The landscape of the reconstruction is in agreement with experimental results [73, 74, 75, 76, 77, 78, 79] and previous DFT calculations [80]. Similar patterns were also seen for the twisted homobilayers [81, 33]. Like previous studies [82, 29, 83, 46, 84, 85, 86], we also decompose the reconstruction into in-plane (IP) and out-of-plane (OOP) displacements. Figures 3(c) and 3(d) show the displacements of the Se atoms on the surface of MoSe2 and WSe2, where the IP displacements are denoted by the white arrows and the OOP displacements are shown in colors. The IP displacements of the Se atoms on the surface of MoSe2 show a clockwise rotation around the HMX site and a counterclockwise rotation around the HXX site [Fig. 3(c)]. The IP displacements reach the maximum (about 0.37 Å for $H$-1.7) in the region between the HMX and HXX sites. The chiral atomic displacements are also seen for the Se atoms on the surface of WSe2 [Fig. 3(d)]. However, they are in opposite directions to those for MoSe2. This trend is consistent with previous studies [87, 88].
@@
For the OOP displacement, the surface Se atoms exhibit fluctuations up to about 5 Å and 4.5 Å for MoSe2 and WSe2, respectively. The fluctuations result in an apex in the HXX region and a basin at the HMM site. The interlayer distance has the largest value at HXX and the smallest value at HMX [Fig. 3(e)], which shows the same trend as the untwisted bilayers [see Fig. 8(a) in Appendix A]. The twisted structures near R-stacking show the same trend in the layer distance ( Appendix C). We quantify the fluctuation in the OOP reconstruction by defining $\delta_{z}=OPP_{max}-OPP_{mini}$, which is shown in Fig. 3(f). It increases monotonically as the twist angle decreases from 21.79 to near 3.15. Then it increases quickly as the angle further decreases. This feature is related to different responses of the layer distances at the three high-symmetry sites to the twist. The layer distances in HMM and HMX regions roughly decrease as the twist angle decreases. In contrast, it increases in the HXX region [Fig. 3(f)]. The results for moiré lattices in $R$-$\theta$ show similar trends in the lattice reconstruction (Appendix C).
@@ III.3 Electronic bands
Figure 4 shows the band structure of MoSe2/WSe2 moiré lattices with the twist angle decreases from 6.01 to 1.7. Here, we focus on the valence bands, among which the highest six bands are labeled by numbers in Fig. 4(d). One can see an almost dispersionless band (band-2) around -0.66 eV at the angle of 3.89, which is 3.7 meV below the valence band maximum at K (VBMK). Its bandwidth is about 4.6 meV and reduces to only 0.7 meV at the angle of 3.15. Then, band-2 moves to be higher than VBMK and becomes increasingly flat. Note that band-3 and band-6 (guided by the red arrows) also move upward quickly as the twist angle decreases. As a result, there are three flat bands for $H$-2.13. Much more flat bands are obtained as the twist angle further decreases down to 1.7. In Fig. 4(j, k), we analyze the wavefunction of the band at $\Gamma$ and K for $H$-3.89 [labeled by the diamonds in Fig. 4(d)], respectively. One can see that the flat band at $\Gamma$ has contributions from both MoSe2 and WSe2 and is mainly localized at HMX. In contrast, the band at K is solely contributed by WSe2. Orbital projections reveal that the flat band at $\Gamma$ is contributed by Se-$p_{z}$ and Mo/W-$d_{z^{2}}$ orbitals. Whereas the band at K is contributed by Se-$p_{x}$/$p_{y}$ and W-$d_{xy}/d_{x^{2}-y^{2}}$ orbitals (see Appendix B Fig. 9). This situation is the same as the untwisted bilayer [89]. Therefore, the flat band (band-2) in Fig. 4(d) is likely formed by the $\Gamma$ valley and the dispersive band (band-1) originates from the K valley, which is further supported by the unfolded bands discussed below. Moreover, one can also deduce that the $\Gamma$-valley flat band can be modulated by both the size of the moiré lattice and interlayer coupling. In contrast, one can expect that the interlayer coupling has a minor effect on the K-valley band, which forms a hexagonal lattice for $H$-3.89.
@@
Figure 4: Electronic structures of twisted MoSe2/WSe2 heterobilayers. (a)-(i) The valence bands of H-$\theta$ ($\theta$ = 6.01, 5.09, 4.41, 3.89, 3.15, 2.88, 2.65, 2.13, and 1.7). The Fermi level is shifted to the middle of the band gap. The red double-headed arrow marks the bandgap between the first and the second flat bands. The leftward arrow is used to track the evolution of band-6 of H-3.89 as the twist angle. The green trajectory is used to mark the K valley. (j, k) The distribution of wave functions of the $\Gamma$ and K valleys (indicated by the red and black rectangles) for H-3.89 in (d). The upper panel and lower panel represent the top view and side view, respectively.
@@
From Fig. 4, one can also see how the flat bands evolve as the twist angle changes. In Fig. 4(a, b), we marked a minigap at K by a circle. The formation of this gap is due to the interaction between the $\Gamma$ valley and its folded band. As the twist angle decreases, the BZ of the moiré lattice becomes smaller and smaller. Therefore, the $k$-point for the gap opening gets close to $\Gamma$. Meanwhile, the interlayer couplings are enhanced due to the reduced interlayer distance at HMX, which yields an increased minigap and pushes band-2 to high energies. These two facts lead to the quick flattening and a dramatic shift of the $\Gamma$-valley band. In contrast, the K-valley band (in green) is relatively less affected. Because it is modulated by the weak moiré potential, the gap openings at $\Gamma$ and M increase much slower than those for the $\Gamma$-valley flat band. Similar behavior is also seen for the moiré lattices in $R$-$\theta$ (see Appendix C).
@@
Figure 5: Unfoled bands for MoSe2/WSe2 moiré lattices near H-stacking. The sizes of the highest minigaps in both the $\Gamma$ and K valleys are shown and guided by the red arrows. The energy difference between the highest $\Gamma$-valley and K-valley flat bands is also shown (guided by the black arrows).
@@
We further obtained the unfolded band structure by projecting the wavefunctions onto the $k$-points in the BZ of the primitive cell of WSe2, for which the results are shown in Fig. 5. The reason is that WSe2 contributes the highest K-valley flat band [see Fig. 9(a) in Appendix B]. Indeed, one can see that for $H$-4.41 and $H$-3.89 the highest $\Gamma$-valley flat band is lower than VBMK. Then it shifts to about 19 meV higher than VBMK for $H$-3.15. Therefore, one can again confirm that band-2 shown in Fig. 4(d) originates from the $\Gamma$ valley. Likewise, by inspecting the energy differences between the bands at $\Gamma$ and K, one can also deduce that the three dispersionless bands (in numbers 1 - 3) shown in Fig. 4(h) are $\Gamma$-valley flat bands. Moreover, an apparent K-valley flat band will not appear until the angle is decreased to 2.13, which is smaller than that for the $\Gamma$ valley. This difference can be understood since the K-valley flat band is mainly caused by the long wavelength moiré potential.

### v2

@@ Abstract
Moiré lattices have served as the ideal quantum simulation platform for exploring novel physics due to the flat electronic bands resulting from the long wavelength moiré potentials. However, the large sizes of this type of system challenge the first-principles methods for full calculations of their electronic structures, thus bringing difficulties in understanding the nature and evolution of the flat bands. In this study, we investigate the electronic structures of moiré patterns of MoSe2/WSe2 by combining ab initio and machine learning methods. We find that a flat band with a bandwidth of about 5 meV emerges below the valence band edge at the K point for the H-stacking at a twist angle of 3.89 without spin-orbit coupling effect. Then, it shifts dramatically as the twist angle decreases and becomes about 20 meV higher than the valence band maximum for the twist angle of 3.15. Multiple ultra-flat bands emerge as the twist angle is reduced to 1.7. The spin-orbit coupling leads to a giant spin splitting comparable to that observed in the untwisted system (about 0.45 eV) and is nearly independent of twisting and stacking. As a result, the K-valley flat band remains the valence band maximum with the inclusion of spin-orbit coupling. Band unfolding reveals that the ultra-flat bands formed by the $\Gamma$ and K valleys show distinct behaviors. The $\Gamma$-valley flat bands are sensitive to the interlayer coupling, thus experiencing dramatic changes as the twist angle decreases. In contrast, the K-valley flat band, which shows a weak dependence on the interlayer coupling, is mainly modulated by structural reconstruction. Therefore, a relatively small angle (2.13) is required to generate the K-valley flat band, which experiences a transition from the honeycomb to the triangular lattice as the twist angle decreases.
@@ I INTRODUCTION
The tbTMDs can exhibit much richer geometric and electronic structures than tBLG [50, 51]. For instance, they can be obtained by twisting the TMD bilayers either in R-stacking or in H-stacking. Namely, their structure is dependent on the relative orientation of the two layers. As a result, they show interesting geometric patterns such as honeycomb and triangular domain wall networks depending on the stacking order. In addition, the TMD bilayers possess two different groups of valleys, i.e., the $\Gamma$ and K valleys. The flat bands originating from these valleys allow for exploring various physical models. For example, the $\Gamma$-valley flat bands have been proposed to realize the honeycomb and kagome lattice models [47], while the K-valley flat bands have been used to study the triangular Hubbard model [40, 52, 53] and the Kane-Mele model [41]. Moreover, the two valleys are energetically close. This feature along with the stacking order and twist angle can give rise to complicated electronic structures for tbTMDs. For instance, for the twisted bilayer of WSe2 with an angle of 5.1 (a twist away from R-stacking) only the $\Gamma$-valley flat bands were observed [55, 56]. As the twist angle decreases to 3, both the $\Gamma$- and K-valley flat bands were observed [45]. For the twists near H-stacking, $\Gamma$-valley flat bands were observed for WSe2 at angles of 57.4 and 57.5 [57, 45]. When the angle is slightly increased to 57.72, the K-valley flat band appears and becomes 129 meV higher than the $\Gamma$-valley one [58].
Among the tbTMDs, twisted heterobilayers, e.g., MoS2/WS2 and MoSe2/WSe2, have emerged as prominent candidates for investigating moiré interlayer excitons [59, 60, 61, 62, 63, 64, 65, 66, 67, 68], which can modulated by the moiré potential. Recently, a scanning tunneling microscope (STM) experiment shows the H-stacking MoSe2/WSe2 moiré system has been reported to exhibit a deep moiré potential exceeding 300 meV for the valence band as the moiré patterns reaching 13 nm. The moiré potential shows a non-monotonic behavior as a function of the moiré wavelength. As the moiré wavelength increases from 6 nm ($\sim$3.15) to 17 nm ($\sim$1.08), a signature of the moiré flat band emerges from the HMX site [69]. However, its valley origin remains to be explored.
@@ II METHODS
We use a DP generated by the DeePMD-kit package to model the interatomic interactions for the MoSe2/WSe2 moiré lattices. The training data were obtained by DFT calculations using the Vienna Ab-initio Simulation Package (VASP) [75]. The projector augmented wave (PAW) method was used to construct pseudopotentials [76]. The energy cutoff for plane waves was 270 eV and the electronic exchange-correlation functional was parametrized by the formalism proposed by Perdew, Burke, and Ernzerhof (PBE) within the generalized gradient approximation [77]. The 2D Brillouin zone (BZ) was sampled using a 5$\times$5 $k$-mesh. The van der Waals forces between layers were corrected using the DFT-D3 method [78]. Then, the LAMMPS program incorporating the DP model was used for structural relaxations [79, 80, 81], for which the threshold force is 10-6 eV/Å. The band structures were obtained using the DeepH-pack package. The atomic-orbital-based $ab$ $initio$ computation program at UStc (ABACUS) [82, 83] was used to generate the training data for DeepH. The pseudo-atomic orbitals used in the ABACUS calculations are Se-$2s2p1d$, Mo-$4s2p2d1f$, and W-$4s2p2d2f$ with a radius cutoff of 10 a.u. and an energy cutoff of 100 Ry. Unfolded bands were obtained using the KPROJ program [84], which projects the wave function of the moiré lattices onto the $k$-points in the BZ of the primitive cell. Figure 1 shows the workflow of our computational scheme.
Figure 1: Workflow of computational scheme.
@@ III.1 DP and DeepH models
For the DP model, the dataset for training was generated by $NVT$ $ab$ $initio$ molecular dynamic (AIMD) simulations using VASP for a few structures, which include 2 $\times$ 2 supercells of the high-symmetry stacking orders shown in Fig. 2, those obtained by a relative shifting of the top layer in the stacking orders, and moiré structures with an angle of 21.79, i.e., $H$-$21.79$ and $R$-$21.79$. In addition, these structures under biaxial strains in the range from -2.5% to 2.5% were also included. For the DeepH model, structures for generating the dataset are 3 $\times$ 3 supercells of the high-symmetry stacking orders, the shifted ones, and the twisted structures with an angle of 13.17. The Hamiltonian hopping and overlapping matrix elements were obtained using ABACUS. Benchmark tests for the DP and DeepH models are shown in Fig. 3. It should be mentioned that we also calculated phonon dispersions for $H$-$13.7$ [Fig. 3(f)] and $R$-$13.7$ [Fig. 3(h)] using the DP model interfaced with LAMMPS and calculated electronic bands for $H$-$21.79$ [Fig. 3(j)] and $R$-$21.79$ [Fig. 3(l)] using DeepH, of which the structures were not included in the training. However, the results from the deep-learning models show good agreements with those from DFT calculations. In addition, we have also performed DeepH calculations for both $H$- and $R$-6.01, whose results are given in Appendix A (see Fig. 11). The band structures derived from DeepH is also in good agreement with those obtained by DFT. These results suggest the high precision of the models.
@@ III.2 Geometric structures
Figure 4(a) shows a rigid moiré structure for $H$-1.7, which exhibits a triangular pattern. The pattern is then changed into a hexagonal domain-wall network after structural relaxation due to geometric reconstruction [see Fig. 4(b)]. The landscape of the reconstruction is in agreement with experimental results [85, 86, 87, 88, 89, 69, 90] and previous DFT calculations [91]. Similar patterns were also seen for the twisted homobilayers [92, 33]. Like previous studies [93, 29, 94, 46, 95, 96, 97], we also decompose the reconstruction into in-plane (IP) and out-of-plane (OOP) displacements. Figures 4(c) and 4(d) show the displacements of the Se atoms on the surface of MoSe2 and WSe2, where the IP displacements are denoted by the white arrows and the OOP displacements are shown in colors. The IP displacements of the Se atoms on the surface of MoSe2 show a clockwise rotation around the HMX site and a counterclockwise rotation around the HXX site [Fig. 4(c)]. The IP displacements reach the maximum (about 0.37 Å for $H$-1.7) in the region between the HMX and HXX sites. The chiral atomic displacements are also seen for the Se atoms on the surface of WSe2 [Fig. 4(d)]. However, they are in opposite directions to those for MoSe2. This trend is consistent with previous studies [98, 99].
@@
For the OOP displacement, the surface Se atoms exhibit fluctuations up to about 5 Å and 4.5 Å for MoSe2 and WSe2, respectively. The fluctuations result in an apex in the HXX region and a basin at the HMM site. The interlayer distance has the largest value at HXX and the smallest value at HMX [Fig. 4(e)], which shows the same trend as the untwisted bilayers [see Fig. 12(a) in Appendix B]. The twisted structures near R-stacking show the same trend in the layer distance. We quantify the fluctuation in the OOP reconstruction by defining $\delta_{z}=OPP_{max}-OPP_{mini}$, which is shown in Fig. 4(f). It increases monotonically as the twist angle decreases from 21.79 to near 3.15. Then it increases quickly as the angle further decreases. This feature is related to different responses of the layer distances at the three high-symmetry sites to the twist. The layer distances in HMM and HMX regions roughly decrease as the twist angle decreases. In contrast, it increases in the HXX region [Fig. 4(f)].
The results for the R-$\theta$ moiré lattices show a similar trend in the lattice reconstruction. As shown in Figs. 5(a) and (b), the relaxed $R$-1.7 shows a pattern of triangular networks, consistent with previous studies and was also seen in the twisted bilayers of other TMD family members  [85, 33, 86, 87, 100, 101, 102, 89, 103, 91, 90, 104, 88, 105]. The displacements of the reconstructed $R$-1.7 are shown in Figs. 5(c) and (d). The IP displacements of the Se atoms on the surface of MoSe2 show a clockwise rotation around the RMX and RXM sites, while a counterclockwise rotation around the RMM site. As expected, the Se atomic displacements of WSe2 are in opposite directions compared to those in MoSe2. The OPP displacements for both the top and bottom layers show the same trend but exhibit different fluctuations, resulting in an apex at the RMM site and a basin at both the RMX and RXM sites. Figure 5(e) shows the interlayer distances for the relaxed $R$-1.7 moiré lattice, which exhibit the largest value at the RMM site. Figure 5(f) shows the changes in the reconstruction ($\delta_{z}$) and the layer distances for high-symmetry sites (RXM, RMX, and RMM) as the twist angle decreases. The behavior of $\delta_{z}$ exhibits a similar trend to that observed in the H-$\theta$ systems.
@@ III.3 Electronic bands
Figure 6 shows the band structure of MoSe2/WSe2 moiré lattices with the twist angle decreases from 6.01 to 1.7. Here, we focus on the valence bands, among which the highest six bands are labeled by numbers in Fig. 6(d). One can see an almost dispersionless band (band-2) around -0.66 eV at the angle of 3.89, which is 3.7 meV below the valence band maximum at K (VBMK). Its bandwidth is about 4.6 meV and reduces to only 0.7 meV at the angle of 3.15. Then, band-2 moves to be higher than VBMK and becomes increasingly flat. Note that band-3 and band-6 (guided by the red arrows) also move upward quickly as the twist angle decreases. As a result, there are three flat bands for $H$-2.13. Much more flat bands are obtained as the twist angle further decreases down to 1.7. In Fig. 6(j, k), we analyze the wavefunction of the band at $\Gamma$ and K for $H$-3.89 [labeled by the diamonds in Fig. 6(d)], respectively. One can see that the flat band at $\Gamma$ has contributions from both MoSe2 and WSe2 and is mainly localized at HMX. In contrast, the band at K is solely contributed by WSe2. Orbital projections reveal that the flat band at $\Gamma$ is contributed by Se-$p_{z}$ and Mo/W-$d_{z^{2}}$ orbitals. Whereas the band at K is contributed by Se-$p_{x}$/$p_{y}$ and W-$d_{xy}/d_{x^{2}-y^{2}}$ orbitals (see Appendix C Fig. 13). This situation is the same as the untwisted bilayer [106]. Therefore, the flat band (band-2) in Fig. 6(d) is likely formed by the $\Gamma$ valley and the dispersive band (band-1) originates from the K valley, which is further supported by the unfolded bands discussed below. Moreover, one can also deduce that the $\Gamma$-valley flat band can be modulated by both the size of the moiré lattice and interlayer coupling. In contrast, one can expect that the interlayer coupling has a minor effect on the K-valley band, which forms a hexagonal lattice for $H$-3.89.
@@
Figure 6: Electronic structures of twisted MoSe2/WSe2 heterobilayers. (a)-(i) The valence bands of H-$\theta$. The Fermi level is shifted to the middle of the band gap. The red double-headed arrow marks the bandgap between the first and the second flat bands. The leftward arrow is used to track the evolution of band-6 of H-3.89 as the twist angle. The green trajectory is used to mark the K valley. (j, k) The distribution of wave functions of the $\Gamma$ and K valleys (indicated by the red and black rectangles) for H-3.89 in (d). The upper panel and lower panel represent the top view and side view, respectively.
@@
From Fig. 6, one can also see how the flat bands evolve as the twist angle changes. In Fig. 6(a, b), we marked a minigap at K by a circle. The formation of this gap is due to the interaction between the $\Gamma$ valley and its folded band. As the twist angle decreases, the BZ of the moiré lattice becomes smaller and smaller. Therefore, the $k$-point for the gap opening gets close to $\Gamma$. Meanwhile, the interlayer couplings are enhanced due to the reduced interlayer distance at HMX, which yields an increased minigap and pushes band-2 to high energies. These two facts lead to the quick flattening and a dramatic shift of the $\Gamma$-valley band. In contrast, the K-valley band (in green) is relatively less affected. Because it is modulated by the weak moiré potential, the gap openings at $\Gamma$ and M increase much slower than those for the $\Gamma$-valley flat band.
Figure 7 shows the evolution of the flat bands in the R-stacking moiré lattice. As the angle decreases from 6.01 to 2.45, the $\Gamma$ valley gradually shifts to a higher energy level than VBMK. A few flat bands emerge as the angle reaches 2.45. However, based on the analysis of the flat bands of the H-stacking moiré lattice, one can expect that these bands originate from the $\Gamma$-valley. We further analyzed the wavefunction of the $\Gamma$-valley flat bands of $R$-3.89 [see Fig. 7(j)], which supports the above expectation. The wavefunction is mainly localized at RMX and RXM sites with interlayer hybridization. However, the RXM sites have contributed more than RMX due to symmetry breaking, which is different from the situation for the twisted homobilayers.
Figure 7: Band structure of R-stacked moiré lattices. (a)-(i) The valence bands of $R$-$\theta$. (j) Charge density distributions of the $\Gamma$ valley (guided by the red diamond) for $R$-3.89 in (d).
There is a difference in the flat bands between the H- and R-stacking moiré structures. The bandwidth of the highest $\Gamma$-valley flat band of the R-stacking moiré patterns is larger than that of the H-stacking at the same angle. For example, the bandwidth of the highest $\Gamma$-valley flat band of R-3.89 is about 9.4 meV, which is double of that for H-3.89 [$\sim$4.6 meV, Fig. 6(d)]. This difference in bandwidth may be related to the localization of high-symmetry stacking configurations. In the H-stacking moiré lattice, the wavefunction of the $\Gamma$-valley flat band is localized only at the HMX site [see Fig. 6(j)]. Whereas the wavefunction of the highest flat band of the R-stacking moiré lattice is localized at both the RMX and RXM sites.
@@
Figure 8: Unfoled bands for MoSe2/WSe2 moiré lattices near H-stacking. The sizes of the highest minigaps in both the $\Gamma$ and K valleys are shown and guided by the green arrows. The energy difference between the highest $\Gamma$-valley and K-valley flat bands is also shown (guided by the white arrows).
@@
We further obtained the unfolded band structure by projecting the wavefunctions onto the $k$-points in the BZ of the primitive cell of WSe2, for which the results are shown in Fig. 8. The reason is that WSe2 contributes the highest K-valley flat band [see Fig. 13(a) in Appendix C]. Indeed, one can see that for $H$-4.41 and $H$-3.89 the highest $\Gamma$-valley flat band is lower than VBMK. Then it shifts to about 19 meV higher than VBMK for $H$-3.15. Therefore, one can again confirm that band-2 shown in Fig. 6(d) originates from the $\Gamma$ valley. Likewise, by inspecting the energy differences between the bands at $\Gamma$ and K, one can also deduce that the three dispersionless bands (in numbers 1 - 3) shown in Fig. 6(h) are $\Gamma$-valley flat bands. Moreover, an apparent K-valley flat band will not appear until the angle is decreased to 2.13, which is smaller than that for the $\Gamma$ valley. This difference can be understood since the K-valley flat band is mainly caused by the structural reconstruction.